# Large-eddy simulations of heat transfer from iced surfaces

### v1

@@ Abstract
Accurate modeling of ice accretion is important for the safe and efficient design of aircraft and wind turbine systems. Heat transfer predictions obtained from the fluid flow solvers are used as input in ice accretion codes. In glaze ice conditions, the freezing rates and resulting ice shapes are highly sensitive to the input values of the heat transfer coefficient. Hence, an accurate prediction of heat transfer on iced airfoils is crucial for correctly predicting the ice accretion process [1, 2]. In this study, we perform conjugate heat transfer (CHT) simulations using wall-modeled large-eddy simulation (WMLES) over surfaces characterized by ice roughness. The results show that WMLES with CHT accurately captures surface temperature distributions and heat fluxes across a range of roughness geometries. For the cases considered, large roughness-to-boundary-layer thickness ratios disrupt outer-layer similarity, leading to substantial errors in estimating equivalent sandgrain roughness when applying traditional empirical models based on surface statistics. The simulations further show that local heat fluxes vary significantly across roughness elements due to the low thermal conductivity of the solid; in particular, roughness crests exhibit reduced fluxes in contrast to slopes and valleys. These effects are not captured in isothermal wall simulations, which overestimate the Stanton number, especially at roughness peaks. By enabling the calculation of Stanton number using heat flux distributions, not directly available in the experiments, the present simulations augment experimental results and highlight the importance of including solid conduction effects for accurately modeling heat transfer over rough, low-conductivity surfaces such as ice.
@@ I Introduction
Two primary types of ice, rime and glaze ice, accrete on aerodynamic surfaces. Rime ice typically occurs at low flight temperatures ($<-20^{\circ}C$), low speeds, low atmospheric liquid water concentration, and small droplet sizes [6]. In rime ice conditions, cooling mechanisms remove all the latent heat of solidification from the impinged supercooled water droplets, freezing them completely upon impact. Under these conditions, the ice shape is determined primarily by the impingement distribution of the droplets and is almost independent of the heat transfer process. In contrast, glaze ice typically forms at near-freezing temperatures, higher speeds, elevated liquid water concentration, and larger droplet sizes. In glaze conditions, not all droplets freeze on impact; a portion of them freeze and the remaining liquid runs back along the airfoil, either freezing further downstream or departing from the surface. This type of ice often appears translucent and can form structures known as horns, which severely degrade aerodynamic performance. In this regime, the rate of ice accretion is highly sensitive to the heat-transfer capacity needed to remove the latent heat of solidification from the droplets.
@@
Numerical ice accretion modeling is typically performed in the following steps: First, a steady-state Reynolds-averaged Navier-Stokes (RANS) solver is used to compute the surface flow field. Next, either an Eulerian or Lagrangian droplet transport model is employed to compute the rate of impact of droplets on the body surface [7, 8, 9]. Information obtained from these models is used as an input to solve a continuity and an energy conservation equation for a water film over the surface, calculating the liquid flow and the freezing rate [10]. Next, the outer mold line of the iced surface is updated, and the domain is remeshed. This process of computing the flow field, droplet collection, and calculating the growth of ice is typically referred to as one shot. When this process is repeated multiple times, it is known as a multishot calculation. In glaze conditions, the freezing rates and resulting ice shapes are highly sensitive to the values of the heat transfer coefficient ($h$) obtained from the simulations and used as input to the ice accretion model [11, 2]. The convective heat transfer coefficient is often obtained from running isothermal wall simulations and then calculating $h$ using $q_{w}=h\times(T_{w}-T_{aw})$, where $q_{w}$ is the wall heat flux, $T_{w}$ is the wall temperature, and $T_{aw}$ is the adiabatic wall temperature [8]. In this approach, the effect of surface roughness on the convective heat‐transfer coefficient $h$ can be incorporated in three ways: (1) by specifying an equivalent sand‐grain roughness height $k_{s}$ to modify the turbulence closure in the RANS equations [12]; (2) by applying empirical correlations developed for icing conditions [13]; and (3) by employing the discrete‐element roughness method to perform a subgrid representation of the roughness elements [14]. Comparisons with experiments have shown that ice accretion codes can overpredict heat transfer coefficients by almost 400% [15]. Results from the Ice Prediction Workshop I revealed a significant disparity in heat transfer coefficients among participants, with peaks ranging from 500 W/m2K to 2500 W/m2K [16]. For these reasons, accurate prediction of heat transfer on iced airfoils is crucial for correctly predicting the ice accretion process [1].
@@
Roughness influences iced surfaces through multiple mechanisms. In the early stages of ice accretion, surface roughness can trip the boundary layer to transition prematurely, increasing both skin friction and heat transfer. As roughness develops, it plays a significant role in enhancing drag and convective heat transfer [17], which affect droplet runback and the dissipation of latent heat. These enhancements are traditionally represented as vertical shifts relative to the logarithmic velocity and temperature profiles of smooth walls [18]. Experiments and direct numerical simulations have shown that enhancement in heat transfer lags behind drag increase [19, 20]. Proportionality between momentum and heat transfer seems to decrease as surface roughness increases. This disparity arises because roughness introduces pressure drag, which significantly impacts momentum transfer but has no direct counterpart in heat transfer. In contrast to drag, that in the fully rough regime is only a function of $k^{+}$, heat transfer remains sensitive to the thermal conductivity.
@@
Roughness-resolved simulations of iced surfaces have been conducted, but these have primarily focused on aerodynamic degradation caused by to ice shape, rather than on heat transfer characteristics [21, 22]. In this article, we present an alternative approach to simulating heat transfer in icing environments through a series of roughness-resolved numerical simulations. We employ a conjugate heat transfer (CHT) solver coupled with wall-modeled large-eddy simulation (WMLES) to model heat transfer in a developing rough-wall boundary layer. The numerical setup replicates the experiments conducted by Ref. 23, who performed heat transfer measurements over a series of flat plates featuring ice-characterized roughness. Heat was generated by mylar plates installed below the testing plate, leading to temperature variations on the rough surfaces that were captured using an infrared (IR) camera installed above it. Here, we compare the temperature distributions and wall heat fluxes obtained from the coupled CHT/WMLES simulations with the experimental measurements. One of the primary goals of this study is to evaluate the ability of WMLES to predict heat transfer on iced surfaces, examining the influence of the wall model in the representation of the temperature boundary layer and the grid resolution required for accurate heat transfer predictions. This study also assesses the role of solid thermal conductivity in modulating heat transfer over rough surfaces and explores its implications for icing simulations. In parallel, we evaluate the performance and limitations of traditional roughness representations based on an equivalent sandgrain roughness. Finally, the simulations presented here serve as a validation case for the CHT/WMLES framework and provide a reference dataset to support the refinement of existing RANS-based models used in icing prediction.
@@
The article is structured as follows. Section II details the CHT/WMLES model, including the governing equations, wall models, and the treatment of different timescales between solid and fluid domains. Section III describes the computational setup. Validation of the numerical model is provided in Section IV. Section V characterizes the influence of surface roughness on the flow and assesses the accuracy of empirical models used for predicting $k_{s}$ for the cases considered. Section VI discusses the heat fluxes obtained from the simulation and the impact of solid conductivity on the Stanton number ($St$) distribution. Finally, Section VII summarizes the study and presents the conclusions.
@@ II Numerical model
II.1 Governing equations
The computational code employed in this study is charLES [24], a low-dissipation, second-order accurate in space, and third-order accurate in time, finite-volume solver. In this work, the low-pass filtered, compressible Navier-Stokes equations for mass, momentum, and total energy,
@@ II.1 Governing equations
$\frac{\partial\overline{\rho}\tilde{u_{i}}}{\partial t}+\frac{\partial\overline{\rho}\tilde{u_{i}}\tilde{u_{i}}}{\partial x_{i}}=-\frac{\partial\overline{p}}{\partial x_{i}}+\frac{\partial\tilde{\tau}_{i,j}}{\partial x_{i}}-\frac{\partial\tilde{\tau}^{sgs}_{i,j}}{\partial x_{i}},$ (2)
$\frac{\partial\overline{E}}{\partial t}+\frac{\partial\tilde{u_{i}}\overline{E}}{\partial x_{i}}+\frac{\partial\tilde{u_{i}}\overline{p}}{\partial x_{i}}=\frac{\partial\tilde{\tau}_{i,j}\tilde{u_{i}}}{\partial x_{i}}-\frac{\partial\tilde{\tau}_{sgs}\tilde{u_{i}}}{\partial x_{i}}+\frac{\partial}{\partial x_{i}}\left(\lambda\frac{\partial\overline{T}}{\partial x_{i}}\right)-\frac{\partial\tilde{Q}^{sgs}_{i}}{\partial x_{i}},$ (3)
are solved, where the operators $\overline{(\cdot)}$ and $\tilde{(\cdot)}$ represent filtered and Favre (density-weighted) filtered quantities, respectively. In these equations, $u_{i}$ stands for the velocity vector, $p$ represents the pressure, $\rho$ is the density, and $\tau_{i,j}$ and $\tau^{sgs}_{i,j}$ represent the viscous and subgrid stresses, respectively. $E$ is the total energy, $T$ is the temperature, and $Q_{i}^{sgs}$ represents the subgrid heat flux. The subgrid terms are closed using the dynamic Smagorinsky model [25]. A constant molecular Prandtl number ($Pr=0.7$) is used to compute the thermal conductivity ($\lambda$), while the subgrid heat flux is computed using a constant turbulent Prandtl number approximation ($Pr_{t}=0.9$). Skew-symmetric operators are employed to conserve kinetic energy in the inviscid, zero Mach number limit, and the scheme also approximately preserves entropy in the inviscid, adiabatic limit [26, 27]. This discretization has been shown to be suitable for coarsely resolved LES of turbulent flows that are especially sensitive to numerical dissipation.
@@
$\frac{\partial\rho_{s}c_{p}T}{\partial t}+\frac{\partial}{\partial x_{j}}\left(\lambda_{s}\frac{\partial T}{\partial x_{j}}\right)=0,$ (4)
@@
where $\rho_{s}$ is the solid density, $c_{p}$ is the heat capacity, $T$ represents the solid temperature, $\lambda_{s}$ denotes the thermal conductivity of the solid, $q_{w,\text{LES}}$ stands for the wall heat flux imposed from the LES calculations, and $q_{\text{radiation}}=\sigma\epsilon(T_{w}^{4}-T_{\infty}^{4})$ is the radiative heat flux, where $\sigma$ is the Stefan–Boltzmann constant, $\epsilon=0.95$ is the emissivity of the ABS plastic, and $T_{\infty}$ is the free-stream temperature. The solution procedures in the fluid and solid domain are tightly coupled; i.e., temperatures and heat-fluxes at the interface are updated every time-step based on the results from the solid and fluid domains, respectively.
II.2 Wall modeling
Due to the high Reynolds number in many engineering applications, resolving the viscous sublayer using LES becomes prohibitively expensive. To obtain accurate results with reasonable computational resources, wall models are used to represent the shear stress and the heat flux at the solid wall [28]. In the present simulations, a laminar wall model is used for the laminar region, and an equilibrium wall model is employed for the turbulent region of the boundary layer. For the laminar region, the shear stress is imposed based on Blasius boundary layer equations [29] as
@@ II.2 Wall modeling
where $U_{\infty}$ is the free-stream velocity. For the equilibrium wall model, we assume that the pressure gradient balances the advective terms, which leads to the constant stress layer approximation,
$\frac{\partial\tau}{\partial n}=\frac{\partial}{\partial n}\left(\mu\frac{\partial u_{t}}{\partial n}-\overline{\rho u_{1}u_{2}}\right)=0,$ (7)
where $u_{t}$ is the velocity component tangential to the wall. The heat flux at the wall is calculated assuming equilibrium assumption in the wall-normal direction:
$\frac{d}{dn}\left(\mu u_{t}\frac{du_{t}}{dn}-u_{t}\overline{\rho u_{1}u_{2}}+\lambda\frac{dT}{dn}-\overline{Tu_{2}}\right)=0.$ (8)
On the interface boundaries between fluid and solid domains, the temperature is imposed based on the solution of the heat equation obtained from the solid side.
II.3 Treatment of solid thermal transients
The substantial differences in timescales between the near-wall turbulent eddies and the heating of the solid necessitate an approach to accelerate the thermal soaking of the solid. In this study, we do so by reducing the thermal inertia of the solid while maintaining a constant Biot number ($Bi$). $Bi$ is a dimensionless number that compares the rate of heat conduction within a solid to the rate of convective heat transfer at its surface:
$Bi=\frac{hL}{\lambda_{s}},$ (9)
where $L$ is a characteristic length scale of the solid, taken here as its thickness. By maintaining a constant Biot number, we prevent the quasi-steady solution from being artificially contaminated [30]. Thermal inertia is initially reduced by a fixed factor, $\beta$. The simulations are started with a thermal inertia set to $\beta\rho c_{p}=10^{-4}\rho c_{p}$. Once a statistically stationary temperature state is reached in the solid, $\beta$ is gradually increased to eliminate unphysical temperature fluctuations caused by the initially low thermal inertia. This gradual increase continues until $\beta$ reaches 1. At $\beta=1$, the surface temperature becomes nearly insensitive to flow fluctuations. A description of the temperature evolution and the effect of $\beta$ is presented in Appendix A.

### v2

# Large eddy simulations of conjugate heat transfer in boundary layers over laser-scanned ice roughness

@@ Abstract
Accurate modeling of ice accretion is important for the safe and efficient design of aircraft and wind turbine systems. Heat transfer predictions obtained from the fluid flow solvers are used as input in ice accretion codes. In glaze ice conditions, the freezing rates and resulting ice shapes are highly sensitive to the input values of the heat transfer coefficient. Hence, an accurate prediction of heat transfer on iced airfoils is crucial for correctly predicting the ice accretion process [1, 2]. In this study, we perform conjugate heat transfer (CHT) simulations using wall-modeled large-eddy simulation (WMLES) over surfaces characterized by ice roughness. The results show that WMLES with CHT accurately captures surface temperature distributions and heat fluxes across a range of roughness geometries. For the cases considered, large roughness-to-boundary-layer thickness ratios disrupt outer-layer similarity, leading to substantial errors in estimating equivalent sandgrain roughness when applying traditional empirical models based on surface statistics. The simulations further show that local heat fluxes vary significantly across roughness elements due to the low thermal conductivity of the solid; in particular, roughness crests exhibit reduced fluxes in contrast to slopes and valleys. Notably, as roughness height increases, the wall heat flux at the crest diminishes, even leading to heat flux reversal in some cases, where thermal energy is transferred from the fluid to the solid. These effects are not captured in isothermal wall simulations, which overestimate the Stanton number, especially at roughness peaks. By enabling the calculation of Stanton number using heat flux distributions, not directly available in the experiments, the present simulations augment experimental results and highlight the importance of including solid conduction effects for accurately modeling heat transfer over rough, low-conductivity surfaces such as ice.
@@ I Introduction
Two primary types of ice, rime and glaze ice, accrete on aerodynamic surfaces. Rime ice typically occurs at low flight temperatures ($20^{\circ}C$), low speeds, low atmospheric liquid water concentration, and small droplet sizes [6]. In rime ice conditions, cooling mechanisms remove all the latent heat of solidification from the impinged supercooled water droplets, freezing them completely upon impact. Under these conditions, the ice shape is determined primarily by the impingement distribution of the droplets and is almost independent of the heat transfer process. In contrast, glaze ice typically forms at near-freezing temperatures, higher speeds, elevated liquid water concentration, and larger droplet sizes. In glaze conditions, not all droplets freeze on impact; a portion of them freeze and the remaining liquid runs back along the airfoil, either freezing further downstream or departing from the surface. This type of ice often appears translucent and can form structures known as horns, which severely degrade aerodynamic performance. In this regime, the rate of ice accretion is highly sensitive to the heat-transfer capacity needed to remove the latent heat of solidification from the droplets.
@@
Numerical ice accretion modeling is typically performed in the following steps: First, a steady-state Reynolds-averaged Navier-Stokes (RANS) solver is used to compute the surface flow field. Next, either an Eulerian or Lagrangian droplet transport model is employed to compute the rate of impact of droplets on the body surface [7, 8, 9]. Information obtained from these models is used as an input to solve a continuity and an energy conservation equation for a water film over the surface, calculating the liquid flow and the freezing rate [10, 11]. Next, the outer mold line of the iced surface is updated, and the domain is remeshed. This process of computing the flow field, droplet collection, and calculating the growth of ice is typically referred to as one shot. When this process is repeated multiple times, it is known as a multishot calculation. In general, ice accretion codes based on a RANS framework are unable to capture certain stochastic features observed in real ice. To address these limitations, several recent models introduce stochasticity directly into the ice accretion framework in an effort to better predict local roughness features [12, 13].
In glaze conditions, the freezing rates and resulting ice shapes are highly sensitive to the values of the heat transfer coefficient ($h$) obtained from the simulations and used as input to the ice accretion model [14, 2]. The convective heat transfer coefficient is often obtained from running isothermal wall simulations and then calculating $h$ using $q_{w}=h\times(T_{w}-T_{aw})$, where $q_{w}$ is the wall heat flux, $T_{w}$ is the wall temperature, and $T_{aw}$ is the adiabatic wall temperature [8]. The effects of surface roughness on the convective heat‐transfer coefficient $h$ are typically incorporated into models via one of three approaches: (1) by specifying an equivalent sand‐grain roughness height $k_{s}$ to modify the turbulence closure in the RANS equations [15, 16]; (2) by applying empirical correlations developed for icing conditions [17]; and (3) by employing the discrete‐element roughness method to perform a subgrid representation of the roughness elements [18]. Comparisons with experiments have shown that ice accretion codes can overpredict heat transfer coefficients by almost 400% [19]. Results from the Ice Prediction Workshop I revealed a significant disparity in heat transfer coefficients among participants, with peaks ranging from 500 W/m2K to 2500 W/m2K [20]. For these reasons, accurate prediction of heat transfer on iced airfoils is crucial for correctly predicting the ice accretion process [1].
@@
Roughness influences iced surfaces through multiple mechanisms. In the early stages of ice accretion, surface roughness can trip the boundary layer to transition prematurely, increasing both skin friction and heat transfer. As roughness develops, it plays a significant role in enhancing drag and convective heat transfer [21], which affect droplet runback and the dissipation of latent heat. These enhancements are traditionally represented as vertical shifts relative to the logarithmic velocity and temperature profiles of smooth walls [22].
@@
Roughness-resolved simulations of iced surfaces have been conducted, but these have primarily focused on aerodynamic degradation caused by to ice shape, rather than on heat transfer characteristics [23, 24]. In this article, we present an alternative approach to simulating heat transfer in icing environments through a series of roughness-resolved numerical simulations. We employ a conjugate heat transfer (CHT) solver coupled with wall-modeled large-eddy simulation (WMLES) to model heat transfer in a developing rough-wall boundary layer. The numerical setup replicates the experiments conducted by Ref. 25, who performed heat transfer measurements over a series of flat plates featuring ice-characterized roughness. Heat was generated by Mylar plates installed below the testing plate, leading to temperature variations on the rough surfaces that were captured using an infrared (IR) camera installed above it. Here, we compare the temperature distributions and wall heat fluxes obtained from the coupled CHT/WMLES simulations with the experimental measurements. One of the primary goals of this study is to evaluate the ability of WMLES to predict heat transfer on iced surfaces, examining the influence of the wall model in the representation of the temperature boundary layer and the grid resolution required for accurate heat transfer predictions. This study also assesses the role of solid thermal conductivity in modulating heat transfer over rough surfaces and explores its implications for icing simulations. In parallel, we evaluate the performance and limitations of traditional roughness representations based on an equivalent sandgrain roughness. Finally, the simulations presented here serve as a validation case for the CHT/WMLES framework and provide a reference dataset to support the refinement of existing RANS-based models used in icing prediction.
@@
The article is structured as follows. Section II details the CHT/WMLES model, including the governing equations, wall models, and the treatment of different timescales between solid and fluid domains. Section III described the experimental data used to validate the numerical model and Section IV describes the computational setup. Validation of the numerical model is provided in Section V. Section VI characterizes the influence of surface roughness on the flow and assesses the accuracy of empirical models used for predicting $k_{s}$ for the cases considered. Section VII discusses the heat fluxes obtained from the simulation and the impact of solid conductivity on the Stanton number ($St$) distribution. Finally, Section VIII summarizes the study and presents the conclusions.
@@ II Numerical model
II.1 Considerations for heat transfer modeling
Before describing the specific numerical approach, we first assess the necessity of using conjugate heat transfer for the present simulations. In order to do so, the dimensionless Biot number ($Bi$) is considered. The Biot number compares the internal thermal resistance due to conduction within a solid to the external resistance due to convection at its surface:
$Bi=\frac{hL}{\lambda_{s}},$ (1)
where $\lambda_{s}$ is the thermal conductivity of the solid, and $L$ is a characteristic length scale of the solid.
For $Bi\ll 0.1$, the internal conduction resistance is much smaller than the convective resistance at the surface, resulting in negligible temperature gradients within the solid. In this regime, the solid can be assumed isothermal and detailed CHT modeling is generally not necessary. However, for $Bi\gtrsim 0.1$, temperature gradients within the solid become significant, since the conduction resistance is comparable to or exceeds the convective resistance. In such cases, CHT is required to resolve the temperature distribution within the solid and accurately predict the surface temperature and heat flux.
In the present work, we estimate the Biot number based on the experimental conditions described in Section III, neglecting the effects of radiation (a conservative assumption). From the experiments, $h\approx 50~\text{W}/(\text{m}^{2}{\cdot}\text{K})$, $\lambda_{s}=0.162~\text{W}/(\text{m}{\cdot}\text{K})$, and the representative length scale is taken as the maximum peak-to-trough roughness height, $k_{z}\approx 0.01~\text{m}$. For this conditions, $Bi\approx 3$, indicating substantial temperature gradients within the roughness elements. Therefore, CHT modeling is essential to accurately simulate the heat transfer process over these rough surfaces.
II.2 Governing equations
The computational code employed in this study is charLES [26], a low-dissipation, second-order accurate in space, and fifth-order accurate in time, finite-volume solver. In this work, the low-pass filtered, compressible Navier-Stokes equations for mass, momentum, and total energy,
@@ II.2 Governing equations
$\frac{\partial\overline{\rho}\tilde{u_{i}}}{\partial t}+\frac{\partial\overline{\rho}\tilde{u_{i}}\tilde{u_{i}}}{\partial x_{i}}=-\frac{\partial\overline{p}}{\partial x_{i}}+\frac{\partial\tilde{\tau}_{ij}}{\partial x_{i}}-\frac{\partial\tilde{\tau}^{sgs}_{ij}}{\partial x_{i}},$ (3)
$\frac{\partial\overline{E}}{\partial t}+\frac{\partial\tilde{u_{i}}\overline{E}}{\partial x_{i}}+\frac{\partial\tilde{u_{i}}\overline{p}}{\partial x_{i}}=\frac{\partial\tilde{\tau}_{ij}\tilde{u_{i}}}{\partial x_{i}}-\frac{\partial\tilde{\tau}^{sgs}_{ij}\tilde{u_{i}}}{\partial x_{i}}+\frac{\partial}{\partial x_{i}}\left(\lambda\frac{\partial\overline{T}}{\partial x_{i}}\right)-\frac{\partial\tilde{Q}^{sgs}_{i}}{\partial x_{i}},$ (4)
are solved, where the operators $\overline{(\cdot)}$ and $\tilde{(\cdot)}$ represent filtered and Favre (density-weighted) filtered quantities, respectively. In these equations, $u_{i}$ stands for the velocity vector, $p$ represents the pressure, $\rho$ is the density, and $\tau_{ij}$ and $\tau^{sgs}_{ij}$ represent the viscous and subgrid stresses, respectively. The subgrid terms are closed using the dynamic Smagorinsky model [27], for which eulerian time averaging is used to determine the Smagorinsky coefficient. The stress tensors are calculated as,
$\tau_{ij}=2\mu\left(S_{ij}-\frac{1}{3}\delta_{ij}S_{kk}\right)$ (5)
$\tau_{ij}^{sgs}=-2C_{s}\rho\Delta^{2}|S|\left(S_{ij}-\frac{1}{3}S_{kk}\delta_{ij}\right)+\frac{1}{3}\tau_{kk}^{sgs}\delta_{ij},$ (6)
with
$|S|=\sqrt{2S_{mn}S_{mn}},\qquad S_{ij}=\frac{1}{2}\left(\frac{\partial\tilde{u}_{i}}{\partial x_{j}}+\frac{\partial\tilde{u}_{j}}{\partial x_{i}}\right),\quad\tau_{kk}^{sgs}\approx 0.$ (7)
$E$ stands for the total energy, $T$ is the temperature, and $Q_{i}^{sgs}$ represents the subgrid heat flux. A constant molecular Prandtl number ($Pr=0.7$) is used to compute the thermal conductivity ($\lambda$), while the subgrid heat flux is computed using a constant turbulent Prandtl number approximation ($Pr_{t}=0.9$), as follows,
$Q_{i}^{sgs}=\frac{1}{Pr_{t}}C_{s}\rho\Delta^{2}|S|\frac{\partial\overline{T}}{\partial x_{i}}.$ (8)
Pressure, density, and temperature are related by the ideal gas law. Viscosity and thermal conductivity are modeled as functions of temperature using a power law with a coefficient of 0.75. Skew-symmetric operators are employed to conserve kinetic energy in the inviscid, zero Mach number limit, and the scheme also approximately preserves entropy in the inviscid, adiabatic limit [28, 29]. This discretization has been shown to be suitable for coarsely resolved LES of turbulent flows that are especially sensitive to numerical dissipation.
@@
$\frac{\partial\rho_{s}c_{p,s}T}{\partial t}+\frac{\partial}{\partial x_{j}}\left(\lambda_{s}\frac{\partial T}{\partial x_{j}}\right)=0,$ (9)
@@
where $\rho_{s}$ is the solid density, $c_{p,s}$ is the solid heat capacity, $T$ represents the solid temperature, $\lambda_{s}$ denotes the thermal conductivity of the solid, $q_{w,\text{LES}}$ stands for the wall heat flux imposed from the LES calculations, and $q_{\text{radiation}}=\sigma\epsilon(T_{w}^{4}-T_{\infty}^{4})$ is the radiative heat flux, where $\sigma$ is the Stefan–Boltzmann constant, $\epsilon=0.95$ is the emissivity of the ABS plastic, and $T_{\infty}$ is the free-stream temperature. Although convection is the dominant mode of heat transfer for the cases presented in this article, the radiative heat flux accounts for approximately $18\%$ of the total heat flux. Neglecting radiation leads to an overprediction of the surface temperature. The solution procedures in the fluid and solid domain are tightly coupled; i.e., temperatures and heat-fluxes at the interface are updated every time-step based on the results from the solid and fluid domains, respectively.
II.3 Wall modeling
Due to the high Reynolds number in many engineering applications, resolving the viscous sublayer using LES becomes prohibitively expensive. To obtain accurate results with reasonable computational resources, wall models are used to represent the shear stress and the heat flux at the solid wall [30]. In the laminar region, the shear stress is imposed based on Blasius boundary layer equations [31] as
@@ II.3 Wall modeling
where $U_{\infty}$ is the free-stream velocity, and the heat flux is imposed assuming equilibrium in the wall-normal direction. For the turbulent region, an equilibrium wall model is employed. We assume that the pressure gradient balances the advective terms, which leads to the constant stress layer approximation,
$\frac{\partial\tau}{\partial n}=\frac{\partial}{\partial n}\left((\mu+\mu_{t\text{,wm}})\frac{\partial\tilde{u}_{t}}{\partial n}\right)=0,$ (12)
where $\tilde{u}_{t}$ and $\tilde{u}_{n}$ are the velocity components tangential and normal to the wall, respectively, and $\mu_{t\text{,wm}}$ is the wall-model eddy viscosity. In this study, we use an algebraic approximation of the wall model derived from the integration of Eq. (12). The heat flux at the wall is calculated assuming equilibrium in the wall-normal direction:
$\frac{d}{dn}\left((\mu+\mu_{t\text{,wm}})\tilde{u}_{t}\frac{d\tilde{u}_{t}}{dn}+\left(\frac{\mu}{Pr}+\frac{\mu_{t\text{,wm}}}{Pr_{t}}\right)\frac{d\overline{T}}{dn}\right)=0.$ (13)
For more detail on the wall model, the reader is referred to [32]. On the interface boundaries between fluid and solid domains, the temperature is imposed based on the solution of the heat equation obtained from the solid side.
II.4 Treatment of solid thermal transients
There are substantial differences in the timescales for heat transfer in the solid and the fluid. While the characteristic thermal diffusion time within the solid is given by $t_{s}=\rho_{s}c_{p,s}L^{2}/\lambda_{s}\approx 1300~\mathrm{s}$, the characteristic time associated with turbulent transport in the boundary layer at the end of the domain is $t_{\text{fluid}}=\delta_{99}/u_{\tau}\approx 0.1~\mathrm{s}$. This large disparity indicates that temperature variations in the solid occur much more slowly than in the fluid. Consequently, it is necessary to employ an approach to accelerate the thermal soaking of the solid. In this study, we employ the equalized timescale method [33], which accelerates thermal soaking by reducing the thermal inertia of the solid. By modifying only the thermal inertia, the Biot number remains unchanged, ensuring that the quasi-steady temperature distribution is unaffected [34]. Thermal inertia is initially reduced by a fixed factor, $\beta$. The simulations are started with a thermal inertia set to $\beta\rho c_{p}=10^{-4}\rho c_{p}$. Once a statistically stationary temperature state is reached in the solid, $\beta$ is gradually increased to eliminate unphysical temperature fluctuations caused by the initially low thermal inertia. This gradual increase continues until $\beta$ reaches 1. At $\beta=1$, the surface temperature becomes nearly insensitive to flow fluctuations. A description of the temperature evolution and the effect of $\beta$ is presented in Appendix A.